# Dynamics of Toxicity in Political Podcasts

### v1

@@ 5.3. Figurative language
Figure 6. Metaphors (left), and, Hyperboles (right) across segments. The numbers are percentages (%).
5.4. Empath features
Empath is a metric which is frequently used in text analytics to assess the emotional content of the text. We also use empath as a metric to understand the underlying emotions and behaviourial patterns latent in the conversation chains. Since our inputs are conversation chains and not flat text, we finetune the DistilBERT (Sanh et al., 2020) model over the dataset presented in (Rashkin et al., 2019) to infer the 32 empath features from these chains. In Figure 7 we show the top eight most frequent emotions that have occurred at least 5% of times within a segment. We observe considerably high intensity for the following empath features – ‘angry’, ‘furious’, ‘annoyed’, ‘disguted’ in the anchor segment. On the other hand, there is a reduced intensity of the following empath features – ‘surprised’, ‘hopeful’, ‘afraid’, ‘anticipating’ in the anchor segment. Thus toxic conversations are laden with aggression resulting in the loss of decorum in the conversation.
@@ 5.4. Empath features
Figure 7. The plot on the left presents the Empath features; those which increased for anchor segment. The plot on the right illustrates the ones with decreased value. In each of these plots, anchor point is ‘0’; left & right semicircles represent previous & next segments respectively.
Figure 8. Examples: Since it is not feasible to illustrate all segments, one among previous and next segments are shown along with anchor segment from the toxic conversation chain. Toxic texts in the anchor segment are marked in red. Note: start and end times are in seconds.
5.5. Keywords
In order to identify and compare the keywords present in the segments we extract the top ten toxic key-phrases from each segment in every conversation chain using KeyBERT (Grootendorst, 2020). Each phrase is limited to a maximum of five-grams. The word embeddings for KeyBERT algorithm are obtained from the Detoxify (Hanu and Unitary team, 2020) library. [^10] https://huggingface.co/unitary/toxic-bert To ensure diversity in our extracted key-phrases, we use the maximal margin relevance utility of the KeyBERT algorithm enforcing a diversity score of 0.75 . In addition, we eliminate stop words & punctuations and only retain strings containing characters from the English alphabet. We combine all the keywords obtained from the preceding ten segments and represent them as a word cloud; similarly, we combine all the keywords obtained from the following ten segments and represent them as another word cloud. We also obtain the word cloud for the anchor segment separately. These word clouds in series are illustrated in Figure 9. We observe that most of the words are linked to political and controversial themes like ‘right’, ‘republican’, ‘democrat’, ‘women’, ‘biden’, ‘american’ & ‘liberal’ which naturally follows from the choice of our dataset. The anchor segment has a number of toxic keywords including ‘idiot’, ‘stupid’, ‘f*cking/f*ck’, ‘sh*t’, ‘a*s’ & ‘moron’. The high similarity between the previous and next word clouds indicate that the anchor segment introduces a disruption in the flow of the main conversation which rewinds back to normal only at the end of the anchor segment. Finally, in both the previous and the next word clouds, words like ‘want’, ‘know’, ‘people’ & ‘yeah’ appear which reflect an expression of demand. Thus hostility in the speech of a particular speaker seem to be fueled by words of demand from either the anchor or other speakers/participants in podcast conversations. .
Figure 9. Word clouds for previous, anchor and next segments.
5.6. Topical shifts
The keyword analysis in the previous section indicates that there is a significant change in the conversation content during the transition from the previous to the anchor segment. This hints to the fact that there is a possible topical/thematic shift during such a transition. In order to establish this we perform topic modeling using the BERTopic (Grootendorst, 2022) model. We extract the topics considering the previous ten aggregated segments a one document, the next ten aggregated segments as a second document and the anchor segment as the third document. We set the number of topics to three and report the top ten most representative words for that topic which has the highest probability of association with a document. The results are noted in Table 1, which reveal significant shifts in thematic focus and toxicity level during the transition from the previous to the anchor segment and the anchor segment to the next segment. Precisely the anchor topic is highly toxic while the preceding and the following topics are more related to demands thus offering insights into the progression and contextual drivers of toxic conversations.
@@ 5.6. Topical shifts
Anchor bitch, stupid, son, fuck, shit, fucking, idiot, damn, shut, guy
@@
6. Change points in toxic chains
While linguistic properties studied in the previous section enables us to characterize the conversation chains, it is not possible to formulate actionable insights for designing effective interventions. In order to achieve this goal one has to acknowledge that toxic conversation chains are inherently dynamic, with evolving interactions that may involve multiple participants. Understanding the trajectory of these conversations requires identifying key points where significant shifts occur. Automatic change point detection can enable researchers to: (i) Isolate key moments: Automatically identify segments of conversations where toxicity sharply increases/decreases or new topics are introduced; and, (ii) Understand contextual triggers: Track down the patterns latent in events or statements that precede toxicity spikes or topic changes and thus reveal actionable insights for designing intervention strategies.
6.1. Automatic change point detection
We utilize standard change point detection (CPD) algorithms to automatically identify change points within conversation chains. Specifically, we consider various search methods – Pelt, KernelCPD, Window, BottomUp & Binseg alongside different cost functions including rbf, cosine, l2 & linear. All implementations are done through ruptures (Truong et al., 2020) library. [^11] https://centre-borelli.github.io/ruptures-docs/ According to the documentation, twelve combinations over these search methods and cost functions are possible; however only rbf as the cost function coupled with Pelt, KernelCPD, BottomUp & Binseg search methods successfully identifies at least one change point. As a result, we conduct our experiments and identify change points using these four algorithms.
6.2. Manual annotation of change points
To evaluate the predictive performance of the baseline CPD algorithms we conduct a manual annotation exercise of the change points. We consider as many as top 100 toxic conversation chains based on the toxicity scores of the anchor segments. We then get all the change points in each of these chains annotated by three annotators. All the three annotators are experts in hate speech analysis research. To ensure consistency and reliability in the annotation process, we ask the annotators to base their judgments on the following factors – (a) tone: shifts in sentiment or intensity, such as transitions from neutral or mildly toxic statements to overt hostility, (b) topical shift: the emergence of new discussion themes or the cessation of previously dominant topics and (c) change in toxicity: escalations or reductions in the degree of harmful language or expressions. Further since this is a sensitive task we instruct the annotators to (a) maintain confidentiality and handle the conversational data responsibly, (b) focus on objective evaluation without imposing personal biases and (c) annotate at most 15 data points per day to ensure annotation quality and mental well being of annotators. We appropriately remunerate the annotators with Amazon gift vouchers.
Overall, the annotators marked a median of 3 change points for most of the chains with the minimum and maximum being 1 and 7 respectively. The total number of change points marked by them across all the 100 chains is 984. We believe that this is a unique and a first-of-its-kind dataset that can be used in future to train/evaluate other predictive algorithms.
6.3. Performance evaluation
Evaluation metrics: In order to measure the correspondence between the manually annotated change points and those generated by the CPD algorithms we use four standard metrics – Hausdorff distance, [^12] https://en.wikipedia.org/wiki/Hausdorff_distance rand index, precision, and recall. While calculating the metrics, if a particular change point is marked by a majority of the annotators it is considered as a valid change point. While the total number of unique points annotated by all 3 annotators is 615, after the majority voting we arrive at 257 points. The predictions of the algorithms are evaluated against these majority voted change points for evaluation.
Key results: The main results are noted in Table 2. We find that the KernelCPD method stands out as the best-performing algorithm across multiple metrics, particularly in terms of Hausdorff distance, rand index, and precision, where it consistently outperforms the other methods. It achieves the highest precision and a high recall value with the margin of two, making it very reliable in detecting change points with both high accuracy and minimal error. Pelt performs decently in most metrics, but lags behind KernelCPD in terms of Hausdorff distance, rand index and precision, indicating that it may miss some important change points. BottomUp provides moderate performance but is generally less reliable than KernelCPD, particularly in terms of precision. Finally, Binseg shows poor results across all metrics, suggesting that it is the least effective algorithm for detecting change points in toxic conversation chains. We present two representative examples of the change points across two different chains in Figure 10.
Metrics
Aggregation
Pelt
KernelCPD
BottomUp
Binseg
avg 6.89 4.66 6.89 7.51
Hausdorff med 7 4 7 7
avg 0.75 0.84 0.72 0.53
rand index med 0.76 0.86 0.74 0.54
avg 0.3 0.53 0.25 0.15
1 med 0.33 0.5 0.25 0
avg 0.37 0.59 0.34 0.22
2 med 0.33 0.57 0.29 0
avg 0.45 0.54 0.35 0.3
precision 4 med 0.4 0.5 0.33 0.33
avg 0.62 0.76 0.72 0.15
1 med 0.6 1 0.75 0
avg 0.74 0.85 0.93 0.22
2 med 0.75 1 1 0
avg 0.88 0.79 0.97 0.29
recall 4 med 1 1 1 0.25
Table 2. Performance comparison of the CPD algorithms across different metrics and two different ways of aggregating the metrics (mean and median). The number 1, 2 and 4 linked with precision and recall are the margins of errors allowed. The best results are highlighted.
Figure 10. Plots representing two samples comparing human annotation with KernelCPD method’s detected change points. Correctly predicted points are marked with a red hexagon and incorrect predictions are marked with red cross.
7. Discussion
Social media has been extensively studied with respect to intervention and demarcation strategies (Jha et al., 2024; Kruzan et al., 2022; Yimam et al., 2024). Rigorous research projects have lead to the development of various strategies like counter speech (Mathew et al., 2019), text detoxification (Dementieva et al., 2024), and meme intervention (Jha et al., 2024), among many others, including some controversial strategies (Leerssen, 2023). Podcasts are comparatively different, since the interaction of the audience with guests/hosts is nearly negligible. Also, unlike social media content, they are generally characterized by long conversations and figuring out exact toxic segments in real time is difficult. Thus, for current scenarios, intervention is limited to only beeping or stripping away such contents which is not real time and is also often ignored by the podcast owners and streaming platforms due to the lack of automation strategies. Further, unlike the access of social media contents, podcasts can easily be made available in offline mode and can create a much bigger indirect impact. In addition, the toxic clips can be spread on social media without any further context thus spicing up the whole thing and leading to even worse polarization.
In this paper we introduced a dataset of transcribed political podcasts and formulated toxic conversation chains from the transcribed text. Firstly, we make the surprising observation on the high prevalence of toxicity (detected with high confidence) in podcasts which reach tens of millions of users. Next, we made various interesting observations regarding the linguistic characteristics of these chains. We noted that the anchor segment corresponding to the most toxic part of the chain has a lot of repetition, is less well-formed and has emotions and sentiments related to anger and fury. Expressions of demand in conversations can quickly lead the discourse to higher levels of toxicity. Subsequently, we manually annotate 100 toxic chains with change points for training and evaluation of automatic methods. We plan to release this dataset upon acceptance. Based on this ground-truth dataset we established that CPD algorithms can be effectively used to monitor the toxicity levels of conversation chains in real time.
The CPD algorithm as demonstrated here can be used to automatically monitor the toxicity levels in real time. Such real time monitoring can be very useful in controlling toxicity by making the participating hosts/guests aware of the rising toxicity in their conversation through a recommendation/alert system. Based on adherence to such recommendations, points, badges and other incentives may be awarded to the participants to promote healthy interaction and deescalate toxicity. In cases of extreme violation automatic termination of the recording of the conversation can also be implemented.
As future work, we plan to extend our study to incorporate audio signals as well accompanied by complete and thorough analysis over different audio features. We also plan to develop intervention strategies for this complex scenario. We also plan to expand our dataset and scalable tooling to other political podcasts to get clear estimates on the prevalence of toxicity in podcasts. Finally, an important aspect that our paper does not answer is the impact of such toxicity on the listeners. Our findings lead us to the question on whether toxicity might become normalized if coming from popular podcast hosts who reach tens of millions of people and the consequences it might have on our political discourse. Some of the tools we developed could be used by social scientists to study and answer such questions.
Ethics statement. Our dataset comprises popular podcasts curated from publicly available RSS feeds, which are widely accessible and listened to by tens of millions of people. These podcasts are analyzed using automated transcription and diarization techniques, and while individual speakers are segmented to enable conversational analysis, their identities are neither inferred nor used in the analysis to ensure privacy. We recognize the potential for biases in various stages of our pipeline, including transcription, diarization, toxicity detection, and change point detection. To mitigate these biases, we employed a rigorous evaluation process, comparing multiple algorithms at each stage and selecting the best-performing ones based on empirical results. While our analysis yielded promising findings, we emphasize the importance of conducting a comprehensive audit of the pipeline to further evaluate its reliability and fairness, particularly in production or high-stakes contexts. Our study adheres to ethical guidelines for research on publicly available data and prioritizes transparency, privacy, and the minimization of harm. We aim to contribute constructively to the discourse on toxicity in podcasts while acknowledging the limitations of our methods and the need for ongoing scrutiny and improvement.

### v2

@@ E.1. Additional textual properties
Figure 11. Token count (left) and type token ratio (right) across segments at 95% confidence intervals.
E.2. Keywords
To identify and compare the keywords present in the segments, we extract the top ten toxic key-phrases from each segment in every conversation chain using KeyBERT (Grootendorst, 2020). Each phrase is limited to a maximum of five-grams. The word embeddings for KeyBERT algorithm are obtained from the Detoxify (Hanu and Unitary team, 2020) library [^24] https://huggingface.co/unitary/toxic-bert. To ensure diversity in our extracted key-phrases, we use the maximal margin relevance utility of the KeyBERT algorithm, enforcing a diversity score of 0.75. In addition, we eliminate stop words & punctuation and only retain strings containing characters from the English alphabet. We combine all the keywords obtained from the preceding ten segments and represent them as a word cloud; similarly, we combine all the keywords obtained from the following ten segments and represent them as another word cloud. We also obtain the word cloud for the anchor segment separately. These word clouds in series are illustrated in Figure 12. We observe that most of the words are linked to political and controversial themes like ‘right’, ‘republican’, ‘democrat’, ‘women’, ‘biden’, ‘american’ & ‘liberal’, which naturally follows from the choice of our dataset. The anchor segment has several toxic keywords including ‘idiot’, ‘stupid’, ‘f*cking/f*ck’, ‘sh*t’, ‘a*s’ & ‘moron’. The high similarity between the previous and next word clouds indicates that the anchor segment introduces a disruption in the flow of the main conversation, which rewinds back to normal only at the end of the anchor segment. Finally, in both the previous and the next word clouds, words like ‘want’, ‘know’, ‘people’ & ‘yeah’ appear, which reflect an expression of demand. Thus, hostility in the speech of a particular speaker seems to be fueled by words of demand from either the anchor or other speakers/participants in podcast conversations.
Figure 12. Word Clouds: previous, anchor and next segments.
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@ E.2. Keywords
Figure 13. Control Group Statistics: Mean at 95% confidence intervals for time coverage, entropy, perplexity, token count & ttr for control group across segments.
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
Figure 14. Left-leaning Statistics: Mean at 95% confidence intervals for time coverage, entropy, perplexity, token count & ttr for left-leaning podcast channels across segments.
@@
Anchor b**ch, stupid, son, f**k, s*it, f**king, idiot, damn, shut, guy
@@
E.3. Topical shifts
The keyword analysis in the previous section indicates that there is a significant change in the conversation content during the transition from the previous to the anchor segment. This hints at the fact that there is a possible topical/thematic shift during such a transition. In order to establish this, we perform topic modeling using the BERTopic (Grootendorst, 2022) model. We extract the topics considering the previous ten aggregated segments as one document, the next ten aggregated segments as a second document and the anchor segment as the third document. We set the number of topics to three and report the top ten most representative words for that topic, which has the highest probability of association with a document. The results are noted in Table 4, which reveals significant shifts in thematic focus and toxicity level during the transition from the previous to the anchor segment and the anchor segment to the next segment. Precisely, the anchor topic is highly toxic while the preceding and the following topics are more related to demands, thus offering insights into the progression and contextual drivers of toxic conversations.
E.4. Number of speakers in toxic chains
We present the number of speaker turns and its distribution over 8,634 toxic chains from right-leaning channels presented as (#speakers: % toxic chain)– (1: 1.26), (2: 8.78), (3: 16.89), (4: 21.35), (5: 19.21), (6: 14.94), (7: 8.92), (8: 4.91), (9: 2.11), (10 and more: 1.63). Notably, only a small percentage of chains, i.e., 1.26% have one speaker within them. Hence, we conclude that the toxic chains represent conversations rather than monologue behaviour.
E.5. Right-leaning control group
Here we cover the analysis of non-toxic chains where we choose such conversations that have the toxicity value of the anchor segment lower than 0.3 (as per the recommendation of Perspective API) for right-leaning channels. We also take care that the randomly selected chains have a distribution in line with the distribution of podcast channels in toxic conversation chains. Further, we exclude such chains that have the previous and next ten segments with toxicity higher than 0.3 to ensure a consistent and fair comparison.
Results are presented in Figure 13. It is evident from the figure that, unlike toxic conversation, non-toxic conversation is more organized & consistent with low randomness.